OpenStudy (anonymous):

Find the number b such that the line y = b divides the region bounded by the curves x = y2 − 1 and the y-axis into 2 regions with equal area. Give your answer correct to 3 decimal places.

OpenStudy (aum):

\(x=y^2-1\) is symmetric about the x-axis because replacing y by -y yields the same equation. So I think the x-axis, or the equation y = 0, divides the regions into two equal areas. So b = 0, unless I am misunderstanding the question.

OpenStudy (anonymous):

Thanks for the explanation, I just graphed it for myself and it seems that b=0 to me as well

OpenStudy (aum):

But if the region they are talking about is bounded by x = y^2 - 1, the y-axis and above the x-axis, then it is a different problem.

OpenStudy (anonymous):

b=0 was incorrect :/ I think it's only the region above the x-axis

OpenStudy (aum):

Then you have to set up two integrals to find the two areas, equate them and solve for b.

OpenStudy (aum):

\[ \int_b^1xdy = \int_0^bxdy \\ \int_b^1(y^2-1)dy = \int_0^b(y^2-1)dy \ \]

OpenStudy (aum):

It is a cubic equation and will have three solutions. Choose the one where 0 < b < 1. Give answer to 3 decimal places.

OpenStudy (anonymous):

Thank you!!!!!

OpenStudy (aum):